### v1

# Retrieve Anyone: A General-purpose Person Re-identification Task with Instructions

Abstract
Human intelligence can retrieve any person according to both visual and language descriptions. However, the current computer vision community studies specific person re-identification (ReID) tasks in different scenarios separately, which limits the applications in the real world. This paper strives to resolve this problem by proposing a new instruct-ReID task that requires the model to retrieve images according to the given image or language instructions. Our instruct-ReID is a more general ReID setting, where existing ReID tasks can be viewed as special cases by designing different instructions. We propose a large-scale OmniReID benchmark and an adaptive triplet loss as a baseline method to facilitate research in this new setting. Experimental results show that the baseline model trained on our OmniReID benchmark can improve +0.5%, +3.3% mAP on Market1501 and CUHK03 for traditional ReID, +2.1%, +0.2%, +15.3% mAP on PRCC, VC-Clothes, LTCC for clothes-changing ReID, +12.5% mAP on COCAS+ real2 for clothestemplate based clothes-changing ReID when using only RGB images, +25.5% mAP on COCAS+ real2 for our newly defined language-instructed ReID. The dataset, model, and code will be available at https://github.com/hwz-zju/Instruct-ReID.
1 Introduction
Identifying individuals exhibiting significant appearance variations from multimodal descriptions is a fundamental aspect of human intelligence with broad applications [29; 27]. To imbue our machine with this capability, person re-identification (ReID) [52; 50] has been introduced to retrieve images of the target person from a vast repository of surveillance videos or images across locations and time [9]. In recent years, significant advancements have been made in developing precise and efficient ReID algorithms and establishing benchmarks covering various scenarios, such as traditional ReID [49; 4; 40], clothes-changing ReID (CC-ReID) [16; 11; 17; 35; 13], clothes template based clothes-changing ReID (CTCC-ReID) [46; 23], and text-to-image ReID (T2I-ReID) [3; 51; 24; 45]. However, focusing solely on one specific scenario, specialized ReID models possess inherent limitations. For instance, customers must employ distinct models to retrieve persons according to languages or clothes templates. To facilitate real-world deployment, there is a pressing need to devise one generic framework capable of re-identifying individuals across all scenarios mentioned above.
This paper proposes a new instruct-ReID task where existing ReID settings can be formulated as special cases. Specifically, the instruct-ReID uses query images and multimodal instructions as the model inputs and requires the model to retrieve the images of the same person in the query image from the gallery following the instructions. Using language or image instructions, models trained on the instructed-ReID task can be specialized to tackle diverse ReID tasks (Fig. 1a). For example, the clothes-changing ReID can be viewed as using the instruction "Ignore clothes" to retrieve images. As another example, clothes template based clothes-changing ReID can utilize a clothes template image as instruction. The proposed instruct-ReID task offers three significant advantages: easy deployment, improved performance, and convenient extension to new ReID tasks. First, it enables cost-effective and convenient deployment in real-world applications. Unlike existing ReID approaches limited to specific tasks, instruct-ReID allows for utilizing a single model for various ReID scenarios, which is more practical in real applications. Second, as all existing ReID tasks can be considered special cases of instruct-ReID, we can unify the training sets of these tasks to exploit the benefits of more data and diverse annotations across different tasks – leading to enhanced performance. Third, instruct-ReID can address a new language-instructed ReID task which requires the model to retrieve persons following language instructions. This setting is very practical in real applications because it enables the customers to retrieve images they are particularly interested in through language descriptions. For example, customers can retrieve a woman wearing a black coat using one of her pictures and language instructions "Wear black coat".
Figure 1: (a) We proposed a new instruct-ReID task that unites various ReID tasks. Traditional ReID: The instruction is "Do not change clothes". The model is expected to retrieve images of the same person without altering image attributes, such as clothing. Clothes-changing ReID: The instruction is "Ignore clothes". The model should retrieve images of the same person even when wearing different outfits. Clothes template based clothes-changing ReID: The instruction is a cropped clothes image. The model should retrieve images of the same person wearing the provided clothing. Language-instructed ReID: The instruction is several sentences describing pedestrian attributes. The model is required to retrieve images of the same person described in the provided sentences. (b) Our proposed method improves the performance of various person ReID tasks by a considerable margin.
To facilitate research in instruct-ReID, we introduce a new benchmark called OmniReID derived from 8 datasets of 4 different ReID tasks. The OmniReID benchmark exhibits two appealing characteristics. First, it emphasizes diversity by incorporating images from various domains, including surveillance scenarios and movie clips. The diversity ensures that the trained models are robust and can effectively handle ReID tasks in various real-world scenarios. Second, the OmniReID benchmark achieves comprehensiveness by offering evaluation datasets to assess various ReID tasks comprehensively, which facilitates evaluating the generalization ability of diverse ReID methods.
Based on OmniReID, we design a baseline method with the proposed adaptive triplet loss for instruct-ReID. The typical triplet loss [34] only defines positive/negative pairs by identities, failing to align with instructions. Therefore, we propose a novel adaptive triplet loss to learn a metric space that preserves identity and instruction similarities. Specifically, we design an adaptive margin between two query-instruction pairs based on instruction similarities to pull features with similar instructions close and push features with different instructions apart. This loss incorporates the instruction information into the features representation and can optimize the model to learn a metric space where similar instructions lead to closer features of query-instance pairs.
In summary, the contributions of this paper are three folds. (1) We propose a new instruct-ReID task, where existing traditional ReID, clothes-changing ReID, clothes template based clothes-changing ReID, and language-instructed ReID can be viewed as special cases. (2) To facilitate research on instruct-ReID, we establish a large-scale and comprehensive OmniReID benchmark consisting 8 publicly available datasets. (3) We propose an adaptive triplet loss to supervise the feature distance of two query-instruction pairs to consider identity and instruction alignments. Our method achieves state-of-the-art results on 7 datasets throughout 4 tasks and on-par results on 1 dataset on traditional ReID tasks. Specifically, our method improves +3.3%, +0.5% mAP on CUHK03, Market1501 for traditional ReID, +2.1%, 15.3% mAP on PRCC, LTCC for clothes-changing ReID when using RGB images only, +12.5% mAP on COCAS+ real2 for clothes template based clothes changing ReID, +25.5% mAP on COCAS+ real2 for our newly defined language-instructed ReID.
2 Related work
Person Re-identification. Person re-identification aims to determine the target images of the same person in the gallery set given the query image. To support the ReID task on all-weather, various tasks are conducted on the scenarios with changing environments, perspectives, and poses [44; 4; 50; 3; 51]. Traditional ReID mainly focuses on dealing with indoor/outdoor and camera pose problems when the target person wears the same clothes. Recently, to extend the application scenarios, clothes-changing ReID (CC-ReID) [46] and clothes template based clothes-changing ReID (CTCC-ReID) [23] are proposed. While CC-ReID forces the model to learn clothe-invariant features [16; 11], CCTC-ReID further extracts clothes template features [23] to retrieve the image of the person wearing the same clothes. In this paper, we introduce a new Instruct-ReID task, which can be viewed as a superset of the existing ReID tasks by incorporating instruction information into identification.
Instruction Tuning. Instruction Tuning was first proposed in the natural language processing (NLP) community to enable large language models (LLM) to execute specific tasks by following natural language instructions. LLM instruction-tuned models, e.g., FLAN-T5 [6], InstuctGPT [31]/ChatGPT [30], can effectively prompt the ability on zero- and few-shot transfer tasks. A few works have borrowed the idea from language to vision. Flamingo [1], BLIP-2 [20], and KOSMOS-1 [15] learning with image-text pairs also show promising generalization on visual understanding tasks. While these LLM-based methods aim to generate convincing language responses following the image or language instructions, we focus on retrieving the correct person following the given instructions by tuning a vision transformer.
Multimodal Retrieval. Multimodel retrieval is a widely used approach that aligns information from multiple modalities to improve model performance of the model and broaden the scope of applications. In multimodel retrieval, unimodal encoders always encode different modalities for retrieval tasks. For instance, CLIP [33], VideoCLIP [42], COOT [10] and MMV [2] utilize contrastive learning for pre-training. Other techniques like HERO [21], Clipbert [19], Vlm [41], bridgeformer [8] and UniVL [28] focus on merging different modalities for retrieval tasks to learn a generic representation. Although there have been numerous studies on multimodal retrieval, most are concentrated on language-vision pretraining or video retrieval, leaving the potential of multimodal retrieval for person ReID largely unexplored. This paper aims to investigate this underexplored area to retrieve anyone with information extracted from multiple modalities.
3 OmniReID Benchmark
To facilitate research on Instruct-ReID, we propose the OmniReID benchmark including a large-scale pretraining dataset based on 8 publicly available datasets with visual and language annotations. The comparison with the existing ReID benchmark is illustrated in Tab. 1.
Table 1: Comparison of training subsets of different ReID datasets.
dataset image ID domain
MSMT17 [40] 30248 1041 indoor/outdoor
CUHK03 [25] 7365 767 indoor
Market1501 [48] 12936 751 outdoor
PRCC [43] 17896 150 indoor
VC-Clothes [39] 9449 256 synthetic
COCAS+ Real1 [23] 34469 2800 indoor/outdoor
OmniReID 121786 5842 indoor/outdoor/synthetic
Protocols. To enable all-purpose person ReID, we collect massive public datasets from various domains and use their training subset as our training subset, including Market1501 [48], MSMT17 [40], CUHK03 [25] for traditional ReID, PRCC [43], VC-Clothes [39], LTCC [32] for clothes-changing ReID, COCAS+ Real1 [23] for clothes template based clothes-changing ReID and language-instructed ReID, forming 121,786 images and 5,842 identities. To fairly compare our method with state-of-the-art methods, the trained models are evaluated on LTCC, PRCC, VC-Clthoes, COCAS+ Real2, Market1501, MSMT17, and CUHK03 test subsets without finetuning. We search the target images with query images and the instructions. Since a query image has multiple target images in the gallery set and CMC (Cumulative Matching Characteristic) curve only reflect the retrieval precision of most similar target images, we also adopt mAP (mean Average Precision) to reflect the overall ranking performance w.r.t. all target images.
Language Annotation Generation. Similar to Text-to-Image ReID datasets, e.g., CUHK-PEDES [24], language annotations in OmniReID are several sentences that describe the visual appearance of pedestrians. We divide our annotation process into pedestrian attribute generation and attribution-to-language transformation.
Pedestrian Attribute Generation. To obtain a comprehensive and varied description of an individual, we employ an extensive collection of attribute words describing a wide range of human visual characteristics. The collection contains 20 descriptions, including full-body clothing, hair color, hairstyle, gender, age, actions, posture, and accessories such as umbrellas or satchels. Professional annotators manually label all the pedestrian attributes. We provide a practical illustration with the attribute collection in Fig. 2(a). By utilizing the instructions on the well-defined attribute combination, models can further enhance their ability to identify the target person. All detailed pedestrian attributes are provided in the supplementary materials.
Attribute-to-Language Transformation. Compared with discrete attribute words, language is more natural for consumers. To this end, we transform these attributes into multiple sentences using the Alpaca-LoRA large language model. Specifically, we ask the Alpaca-LoRA with the following sentences: "Generate sentences to describe a person. The above sentences should contain all the attribute information I gave you in the following." Annotators carefully check generated annotations to ensure the correctness of the language instructions. We provide more language annotations of images in the supplementary materials.
Figure 2: (a) We first generate attributes for a person and then transform attributes into sentences by a large language model. (b) We crop the upper clothes and search them online for clothes templates.
Visual Annotation Generation. Visual annotations are generally images that describe the characteristics of pedestrians. In this paper, we select clothes as the visual annotations because they are viewed as the most significant visual characteristics of pedestrians. To obtain high-quality visual annotations, we first crop the upper clothes from the source images (Clothes Copping) and search for the upper clothes on the internet to get the most corresponding cloth-template images (Clothes-template Crawling) as visual annotations. Since each person wears the same clothes in the traditional ReID dataset, we only annotate the clothes-changing ReID datasets where each person wears multiple clothes to ease the burden of annotations. We detail the process in Fig. 2 (b).
Clothes Cropping. We use a human parsing model, SCHP [22], to generate the segmentation mask of the upper clothes and then crop the corresponding rectangle upper clothes patches from the original images. These bounding boxes of upper clothes are then manually validated.
Clothes-template Crawling. Given all cropped upper clothes from images in OmniReID datasets, we crawl the templates of these clothes from shopping websites [^1] https://world.taobao.com/, https://www.17qcc.com/. The top 40 matching clothes templates are downloaded when we crawl each cropped upper clothes. The one with the highest image quality is manually selected.
Figure 3: The overall architecture of the proposed method. The instruction is fed into the instruction encoder to extract instruction features. The features are then propagated into the editing transformer to capture instruction-edited features. We exploit adaptive triplet loss and identification loss to train the network. For the testing stage, we use the CLS token for retrieval.
4 RAM: Retrieve Anyone Model
The proposed retrieve anyone model is shown in Fig. 3. Our model consists of three parts: instruction generation for different ReID tasks, an editing transformer $\mathcal{E}_{e}$ and an instruction encoder $\mathcal{E}_{i}$ which is the fixed FashionCLIP model [5], which we show better discriminative abilities than CLIP model [33] in the supplementary materials. Given an instruction $\mathbf{T}$ associated with a query image $\mathbf{I}$ (Sec. 4.1), we extract instruction features $\mathbf{F}_{\mathbf{T}}$ by the instruction encoder $\mathcal{E}_{i}$. Then the extracted instruction feature $\mathbf{F}_{\mathbf{T}}$ and the query image tokens are fed into the designed editing transformer to get the features $\mathbf{F}$ edited by institutions (Sec. 4.2). We impose an identification loss $\mathcal{L}_{id}$, which is the classification loss on identities, and an adaptive triplet loss $\mathcal{L}_{atri}$ on $\mathbf{F}$ to supervise the learning of model. The adaptive triplet loss is utilized to pull features with the same identity and similar instructions close and push features with dissimilar instructions apart (Sec. 4.3).
4.1 Instruction Generation
In our proposed instruct-ReID task, the model must retrieve the images that describe the same person following the instructions. By designing different instructions, our instruct-ReID can be specialized to existing ReID tasks, i.e., traditional ReID, clothes-changing ReID, clothes template based clothes-changing ReID, and language-instructed ReID. We show the current instructions below and leave the exploration of better instructions for instruct-ReID to future research.
Traditional ReID: The instruction is the sentence "Do not change clothes". The model is expected to retrieve images of the same person without altering image attributes, such as clothing. ### Query image:{Query image}  ### Instruction: "Do not change clothes."
### Target image:{Output}
Clothes-changing ReID: The instruction is the sentence "Ignore clothes". The model should retrieve images of the same person even when wearing different outfits. ### Query image:{Query image}  ### Instruction: "Ignore clothes."
### Target image:{Output}
Clothes template based clothes-changing ReID: The instruction is a clothes template for a query image while a cropped clothes image for a target image. The model should retrieve images of the same person wearing the provided clothes. We provide more examples for training and test in the supplementary materials. ### Query image:{Query image}  ### Instruction:{Any clothes template image}
### Target image:{Output}
Language-instructed ReID: The instruction is several sentences describing pedestrian attributes. We randomly select the description languages from the person images in gallery and provide to query images as instruction. The model is required to retrieve images of the same person described in the provided sentences. We provide more examples for training and test in the supplementary materials. ### Query image:{Query image}  ### Instruction:{Sentences describing pedestrians}
### Target image:{Output}
4.2 Editing Transformer
The editing transformer consists of $L$ zero-init transformer layers $\mathcal{E}_{e}=\{\mathcal{F}_{1},\mathcal{F}_{2},...,\mathcal{F}_{L}\}$, which can leverage the instruction to edit the feature of query images. Given the $l$-th zero-init transformer, the output feature $\mathbf{F}_{l}$ can be formulated as
$\small\mathbf{F}_{l}=\mathcal{F}_{l}(\mathbf{F}_{l-1},\mathbf{F}_{\mathbf{T}}),$ (1)
where $\mathbf{F}_{\mathbf{T}}=\mathcal{E}_{i}(\mathbf{I})$ is the instruction feature extracted by the instruction encoder $\mathcal{E}_{i}$ and $\mathbf{F}_{l-1}$ is the output feature of $(l\!-\!1)$-th zero-init transformer layer. Here, we define the initial input $\mathcal{F}_{0}$ of the first layer as $\mathbf{F}_{0}=[\mathbf{f}_{0}^{\text{CLS}},\mathbf{f}_{0}^{1},\mathbf{f}_{0}^{2},...,\mathbf{f}_{0}^{N}],$ where $\mathbf{f}^{\text{CLS}}_{0}$ is the appended [CLS] token, $(\mathbf{f}_{0}^{1},\mathbf{f}_{0}^{2},...,\mathbf{f}_{0}^{N})$ are the patch tokens of the query image and $N$ is the number of patches of the query image.
We illustrate the detailed structure of each layer in the editing transformer in Fig. 3 (right). Given the features $\mathbf{F}_{l-1}$ and instruction features $\mathbf{F}_{\mathbf{T}}$, the attention map $\mathbf{M}_{l}$ is defined as
$\small\mathbf{M}_{l}=\left[\text{Softmax}(\mathbf{S}_{l}),g_{l}\times\text{Softmax}(\mathbf{S^{\prime}}_{l})\right],$ (2)
where $g_{l}$ is the gating parameters initialized by zero. Here, $\mathbf{S}_{l}$ is the attention map between queries and keys of input features and $\mathbf{S}^{\prime}_{l}$ is the attention map between queries of input features and keys of instruction features. Mathematically,
$\small\mathbf{S}_{l}=\mathbf{Q}_{l}\mathbf{K}_{l}^{\top}/\sqrt{C},\mathbf{S}_{l}^{\prime}=\mathbf{Q}_{l}\mathbf{K^{\prime}}_{l}^{\top}/\sqrt{C},$ (3)
where a linear projection derives queries and keys, i.e., $\mathbf{Q}_{l}\!=\!\text{Linear}_{q}(\mathbf{F}_{l-1})$, $\mathbf{K}_{l}\!=\!\text{Linear}_{k}(\mathbf{F}_{l-1})$ and $\mathbf{K}^{\prime}_{l}\!=\!\text{Linear}_{k^{\prime}}(\mathbf{F}_{\mathbf{T}})$, respectively. $C$ is the feature dimension of query features. Finally, we calculate the output of the $l$-th layer by
$\small\mathbf{F}_{l}=\text{Linear}_{o}(\mathbf{M}_{l}\left[\mathbf{V}_{l},\mathbf{V}^{\prime}_{l}\right]),$ (4)
where $\text{Linear}_{o}$ is the feed-forward network after the attention layer in each transformer block, $\mathbf{V}_{l}$ and $\mathbf{V}^{\prime}_{l}$ are the values calculated by $\mathbf{V}_{l}=\text{Linear}_{v}(\mathbf{F}_{l-1})$ and $\mathbf{V}^{\prime}_{l}=\text{Linear}_{v^{\prime}}(\mathbf{F}_{\mathbf{T}})$, respectively. We use the [CLS] token in the output feature of $L$-th transformer layer for computing losses and retrieval, i.e., $\mathbf{F}=\mathbf{f}_{L}^{\text{CLS}}$, where $\mathbf{F}_{L}=(\mathbf{f}_{L}^{\text{CLS}},\mathbf{f}_{L}^{1},\mathbf{f}_{L}^{2},...,\mathbf{f}_{L}^{N})$ and $N$ is patch number of query images.
4.3 Adaptive Triplet Loss
Unlike typical triplet loss that defines positive and negative samples solely based on identities, instruct-ReID requires distinguishing images with different instructions for the same identity. Intuitively, an adaptive margin should be set to push or pull samples based on the instruction difference. Let ($\mathbf{F}^{a}_{i}$, $\mathbf{F}^{r_{1}}_{i}$, $\mathbf{F}^{r_{2}}_{i}$) be the $i$-th triplet in the current mini-batch, where $\mathbf{F}^{a}_{i}$ is an anchor sample, $\mathbf{F}^{r_{1}}_{i}$ and $\mathbf{F}^{r_{2}}_{i}$ are reference samples. We propose a novel adaptive triplet loss as
$\small\mathbf{\mathcal{L}}_{atri}=\frac{1}{N_{tri\dagger}}\sum_{i=1}^{N_{tri\dagger}}\left\{\mathbf{Sign}(\beta_{1}-\beta_{2})\left[d(\mathbf{F}^{a}_{i},\mathbf{F}^{r_{1}}_{i})+\left(\beta_{1}-\beta_{2}\right)m-d(\mathbf{F}^{a}_{i},\mathbf{F}^{r_{2}}_{i})\right]\right\}_{+}$ (5)
where $N_{tri\dagger}$ and $m$ denote the number of triplets and a hyper-parameter for the maximal margin, respectively. $d$ is a Euclidean distance function, i.e., $d(\mathbf{F}^{a}_{i},\mathbf{F}^{r}_{i})=\left|\left|\mathbf{F}^{a}_{i}-\mathbf{F}^{r}_{i}\right|\right|_{2}^{2}$. $\beta_{1}$ and $\beta_{2}$ are relatednesses between the anchor image and the corresponding reference image that take into account the identity consistency and the instruction similarity for the adaptive margin. Mathematically, they can be defined as
$\small\beta_{j}=\mathbb{I}\left(y_{a}=y_{r_{j}}\right)\mathbf{Cos}\left\langle\mathbf{F}_{\mathbf{T}}^{a},\mathbf{F}_{\mathbf{T}}^{r_{j}}\right\rangle,$ (6)
where $y_{a}$ and $y_{r_{j}}$ are the identity labels of the anchor image and the reference image, $\mathbb{I}(\cdot)$ is the indicator function, and $j=\{1,2\}$ denotes the index of reference samples.
The concept of adaptive triplet loss can be described by Fig. 4(a). We discuss the adaptive loss in two scenarios. In the first scenario, as shown in Fig. 4(b), the margin is set to zero if the triplet has the same identity and the instructions of the two reference samples are equally similar to the instruction of the anchor sample. This makes the distances between the anchor point and the two reference points the same. Second, as shown in Fig. 4(c), when a significant difference in the instruction similarities exists, the margin distance between the anchor and two references turns to be more than the maximum value $m$, forcing the model to learn discriminative features.
Figure 4: Illustration of adaptive triplet loss. Unlike traditional triplet loss where the margin is fixed, the margin in our adaptive triplet loss is defined by the instruction similarity for the two query-instruction pairs that describe the same person. The features associated with similar instructions are pulled to be close to each other.
5 Experiment
5.1 Implementation Details
Training Settings. To enable all-purpose person ReID, we perform two training scenarios based on the built benchmark OmniReID: 1) Single-task Learning (STL): Every task is trained and tested individually using the corresponding dataset. 2) Multi-task Learning (MTL): To acquire one unified model for all tasks, the model is optimized by joint training of the four ReID tasks with all the training datasets. The trained network is then evaluated for different tasks on various datasets. We provide the details of datasets used in STL and MTL in the supplementary materials.
Training Setups. For the editing transformer, we use the plain ViT-Base with the supervised ImageNet pretraining [38]. The instruction encoder is a fixed FashionCLIP [5]. All images are resized into 256$\times$128 for training and testing. We use the SGD optimizer with a base learning rate of 1e-3 and a weight decay of 5e-4 . We linearly warmup the learning rate from 1e-5 to 1e-3 for the first 1500 iteration steps. For STL scenario, only one V100 GPU is used to train the network with 24000 iterations. For MTL scenario, eight V100 GPUs are used. To deal with multi-task distributed training, following the practice in [37; 7], each GPU is responsible for one dataset during training. Random cropping, flipping, and erasing are used for data augmentation during training. For each training batch, we randomly select 32 identities with 4 image samples for each identity.
5.2 Experimental Results
Clothes-Changing ReID (CC-ReID). As observed in Table 2, RAM outperforms all state-of-the-art methods on LTCC, PRCC and VC-Clothes, showing that the model can effectively extract clothes-invariant features following the "Ignore clothes" instruction. Specifically, on STL, RAM improves the performance of CAL[11], TransReID [12] by +8.9% mAP and +2.9% mAP on LTCC and VC-Clothes, respectively. Using multi-task training further improves the performance of RAM to 56.1% mAP on LTCC, and to reach a new state-of-the-art on PRCC with a 48.0% mAP. While multi-task learning leads to slightly lower performance, our method still achieves a higher Top-1 than TransReID. We conjecture that the performance drop is due to the domain gap between VC-Clothes (Synthetic) and other datasets (Real) and leave it for feature work.
Table 2: The performance of Clothse-Changing ReID of our method and the state-of-the-art methods. Mean average precison (mAP) and Top1 are used to quantify the accuracy.
Methods Modality LTCC PRCC VC-Clothes
mAP Top1 mAP Top1 mAP Top1
HACNN[26] RGB 26.7 60.2 - 21.8 - -
RGA-SC[47] RGB 27.5 65.0 - 42.3 67.4 71.1
PCB[36] RGB 30.6 65.1 38.7 41.8 62.2 62.0
IANet[14] RGB 31.0 63.7 45.9 46.3 - -
CAL[11] RGB 40.8 74.2 - - - -
TransReID[12] RGB - - - 44.2 71.8 72.0
RAM (STL) RGB 49.7 89.0 41.5 41.5 74.7 86.7
RAM(MTL) RGB 56.1 92.9 48.0 47.8 72.0 84.9
Clothes Template Based Clothes-Changing ReID (CTCC-ReID). Our method achieves desirable performance on the CTCC-ReID task in Tabel 3, which shows that a fixed instruction encoder is enough for this tough task. Concretely, when only trained on COCAS+ Real1, RAM outperforms BC-Net [46] and DualBCT-Net [23], both of which learn an independent clothes branch, by +8.6% mAP and +1.2% mAP, respectively. By integrating the knowledge on other instruct ReID tasks during multi-task learning, we are able to further improve the performance of RAM, achieving an mAP of 42.5% and pushing the performance limits on CTCC-ReID.
Table 3: Performance comparison with the state-of-the-art methods on Clothes -Template Clothes-Changing ReID, Language-Instructed ReID, and Traditional ReID. † denotes that the model is trained with multiple datasets.
Methods CTCC-ReID LI-ReID Trad-ReID
COCAS+ Real2 COCAS+ Real2 Market1501 MSMT17 CUHK03
mAP Top1 mAP Top1 mAP Top1 mAP Top1 mAP Top1
Baseline - - 14.9 31.6 - - - - - -
TransReID [12] 5.5 17.5 - - 86.8 94.4 61.0 81.8 - -
Pixel Sampling [35] 2.1 11.6 - - - - - - - -
BC-Net [46] 22.6 36.9 - - - - - - - -
DualBCT-Net [23] 30.0 48.9 - - - - - - - -
SAN [18] - - - - 88.0 96.1 - - 76.4 80.1
HumanBench† [37] - - - - 89.5 - 69.1 - 82.6 -
RAM (STL) 31.2 51.7 31.1 56.9 88.0 99.8 65.6 92.8 85.0 99.1
RAM (MTL)† 42.5 62.9 40.4 60.6 90.0 99.8 69.1 93.7 85.9 99.1
Language-Instructed ReID (LI-ReID). As a new setting in ReID, no previous works have been done to retrieve a person using several sentences as the instruction, therefore, we compare RAM with a straightforward baseline. In the baseline method, only a ViT-Base is trained on COCAS+ Real1 images without utilizing the information from language instruction, leading to poor person re-identification ability. As shown in Table 3, RAM improves the baseline by +16.2% mAP, because RAM can integrate instruction information into identity features. With MTL, RAM achieves extra +9.3% performance gain by using more images and general information in diverse ReID tasks.
Traditional ReID (Trad-ReID). RAM also shows its effectiveness on Traditional ReID tasks in Table 3. Specifically, when trained on a single dataset, compared with SAN [18], RAM achieves comparable performance on Market1501 and +8.6% performance gain on CUHK03. With multi-task training, RAM can further outperform the recent multi-task pretraining SOTA HumanBench [37].
5.3 Ablation Study
Table 4: Ablation study. The performance comparisons between fusion methods, and the comparisons with triplet loss and the proposed adaptive triplet loss in terms of mAP.
Methods CTCC-ReID LI-ReID CC-ReID Trad-ReID Avg.
Real2 Real2 LTCC PRCC VC-Clothes Market1501 MSMT17 CUHK03
RAM (STL) 31.2 31.1 49.7 41.5 74.7 88.0 66.9 85.0 58.5
Late Fusion 31.3 30.2 49.0 41.3 73.1 87.4 65.6 84.5 57.8
Triplet loss 30.1 30.1 49.7 41.5 74.7 88.0 66.9 85.0 58.3
RAM (MTL) 42.5 40.4 56.1 48.0 72.0 90.0 69.1 85.9 63.0
Late Fusion 41.4 37.6 55.6 47.3 71.6 89.6 67.6 85.0 62.0
Triplet loss 42.2 39.7 55.0 47.0 71.0 89.1 66.8 85.0 62.0
Editing Transformer. To verify the effectiveness of the instruction integrating design in the editing transformer, we compare it with Late Fusion, where dual cross-attentions are used between tokens from different models/modalities to fuse information. Results in Table 4 show that adopting Late Fusion leads to -0.7% mAP performance drop in the STL scenario. Consistent results can be observed in MTL, inidicating the effectiveness of the instruction integrating design in the editing transformer.
Adaptive Triplet Loss. Table 4 shows that adaptive triplet loss in MTL outperforms the traditional triplet loss by +1.0% mAP on average, indicating that the proposed loss boosts the model to learn more discriminative features following different instructions. In the STL scenario, for CC-ReID and Trad-ReID, the instructions are fixed sentences leading to the same performance of adaptive/traditional triplet loss. However, in the case of CTCC-ReID and LI-ReID, where instructions vary among samples, using adaptive triplet loss brings about +1.0% mAP performance gain, which demonstrates the effectiveness of adaptive triplet loss in learning both identity and instruction similarity.
6 Visualizaton
We visualize the retrieval results of CTCC-ReID, LI-ReID, CC-ReID and Trad-ReID tasks in Fig. 5. Given a query image, RAM not only retrieves the right person from the gallery but also finds specific target images of the person following the instruction. Concretely, for CTCC-ReID, RAM retrieves images of query persons wearing instructed clothes as shown in the first row. For LI-ReID, RAM effectively parses information from languages such as bag condition (e.g., row 2 person #1) and clothes attribute (e.g., row 2 person #2,3) to find the correct image. For CC-ReID, with the "Ignore clothes" instruction, our method focuses on biometric features and successfully retrieves the person in the case of changing clothes, e.g. the 4th and 5th images of row 3 person #1, which the clothes are different from the query image. For Trad-ReID shown in row 4, "do not change clothes" instructs RAM to pay attention to clothes, a main feature of a person’s image. In this case, RAM retrieves images with the same clothes in the query image.
Figure 5: Illustration of all tasks retrieval results by RAM. We visualize the task-specific instructions on three people for example. Green and red boxes mean true and false matches.
7 Conclusion and Limitations
In this paper, we tackle existing traditional ReID (Trad-ReID), clothes-changing ReID (CC-ReID), clothes template based clothes-changing ReID (CTCC-ReID) and language-instruct ReID (LI-ReID) by proposing a new instruct-ReID task, which holds great potential and broad impact in the security application, scientific research and other fields. For a specific task case, the target person images in the gallery are retrieved by jointly considering the query person image and given instructions. Based on instruct-ReID, we establish a large-scale and comprehensive OmniReID benchmark with the existing publicly available datasets. Furthermore, we design a generic framework RAM for the proposed instruct-ReID with an adaptive triplet loss to consider identity and instruction alignments. One limitation of this work is that instruct-ReID only implements four ReID tasks with task-specific instructions. We hope our OmniReID can facilitate future works such as unified network structure design and multi-task learning methods on a broad variety of retrieval tasks.
Appendix
Appendix A Details of Language Annotation Generation
In this section, we provide more details about the generation of language annotation in OmniReID. To depict pedestrian images with more detailed descriptions, we first manually annotate OmniReID with local description attribute words, including clothes color, accessory, pose, etc. From a human perspective, language descriptions with sentences are much effective in describing person than simply listing attribute. Based on the attribute annotations we collect, we merge them into language annotations to provide a more comprehensive description of the person.
A.1 Pedestrian Attribute Generation
We annotate 20 attributes and 92 specific representation in words for OmniReID, as shown in Table. 5. The attributes are carefully selected considering a wide range of human visual characteristics from the datasets, including full-body clothing, hair color, hairstyle, gender, age, actions, posture, and accessories such as umbrellas or satchels. Though there might be more than one representation for attributes such as coat color and trousers color, only one representation corresponding to the image is selected for annotation. The attributes are annotated by professional annotators in the image level, thus the annotation file will contain more accurate and detailed description. For example, in Fig. 6(a), although the two images are of the same identity, the difference of pedestrian angle is also annotated.
Table 5: Details of attributes from OmniReID that describe person images. We select 20 attributes and 92 specific representation words considering a wide range of human visual appearances in detail.
Attribute representation in words
coat color
"black coat", "blue coat", "gray coat", "green coat", "purple coat", "red coat", "white coat", "yellow coat"
trousers color
"black trousers", "blue trousers", "gray trousers", "green trousers", "purple trousers", "red trousers", "white trousers", "yellow trousers"
coat length
"agnostic length coat", "long sleeve coat", "short sleeve coat", " bareback coat"
trousers length
"shorts trousers", "skirt", "trousers"
gender code
"female", "agnostic gender", "male"
glass style
"without glasses", "with glasses", "with sunglasses"
hair color
"black hair", "agnostic color hair", "white hair", "yellow hair"
hair style
"bald hair", "agnostic style hair", "long hair", "short hair"
bag style
"backpack", "hand bag", "shoulder bag", "waist pack", "trolley", "agnostic style bag", "without bag"
cap style
"with hat", "without hat"
shoes color
"black shoes", "blue shoes", "gray shoes", "green shoes", "purple shoes", "red shoes", "white shoes", "yellow shoes"
shoes style
"boots", "leather shoes", "sandal", "walking shoes"
age
"adult", "child", "old"
person angle
"back", "front", "side"
pose
"lie", "pose agnostic", "sit", "stand", "stoop"
coat style
"business suit", "agnostic style coat", "dress", "jacket", "long coat", "shirt", "sweater", "t-shirt"
glove
"with glove", "agnostic glove", "without glove"
smoking
"smoking", "agnostic smoking", "without smoking"
umbrella
"with umbrella", "without umbrella"
uniform
"chef uniform", "common clothing", "firefighter uniform", "medical uniform", "office uniform", "agnostic uniform", "worker uniform"
A.2 Attribute-to-Language Transformation
We provide some example from our OmniReID that provides sentence descriptions of individuals in images. Compared with discrete attribute words, language is more natural for consumers. To this end, we transform these attributes into multiple sentences using the Alpaca-LoRA large language model. Specifically, we ask the Alpaca-LoRA with the following sentences: "Generate sentences to describe a person. The above sentences should contain all the attribute information I gave you in the following." Annotators carefully check the generated annotations to ensure the correctness of the language instructions. Fig. 6a presents the examples of the same identity and Fig. 6b presents the examples of the transformation with different domains and identities.
Figure 6: We first generate attributes for a person and then transform attributes into sentences by a large language model. (a) The attributes and language guidance of same identity. (b) More instances of attributes-to-language transformation with different domains and identities.
Appendix B The Discriminative Abilities of FashionCLIP and CLIP
In this section, we compare the capacity of fixed FashionCLIP model [5] and CLIP model [33] as instruction encoder to distinguish different instructions. As shown in Fig. 7, the instruction features are extracted from upper clothes images and we calculate the similarity between template upper clothes and cropped upper clothes images. CLIP model is pretrained on classification dataset which is lack of discernment of human details. Take the first row of Template1 for instance, the similarity score between template clothes and corresponding cropped clothes (in green box) is only 0.698, which is even smaller than the second, forth, fifth and sixth cropped clothes images. The counter-intuitive result indicates that the CLIP model is unable to distinguish texture, color and other style characteristics. The FashionCLIP is a CLIP-like model for the fashion industry and fine-tuned on fashion retrieval dataset. The similarity scores of each template between template clothes and corresponding cropped clothes in the second row are obviously exceed others, which indicates that FashionCLIP model [5] shows better discriminative abilities than CLIP model [33].
Figure 7: We extract instruction features respectively by CLIP model or FashionCLIP model and calculate the instruction similarity scores of upper clothes templates and the cropped upper clothes. Green boxes mean the template clothes images is searched on the internet by these corresponding cropped clothes images.
Appendix C Instruction Generation
In our proposed instruct-ReID task, each identity is further split into query and gallery, where query set consists of query person images and clothes templates, and gallery set consists of target person images. In this section, we provide more training and test examples for Clothes template based clothes-changing ReID and Language-instructed ReID scenarios.
C.1 Clothes Template Based Clothes-Changing ReID
In CCTC-ReID scenario, as shown in Fig. 8, the instruction is a clothes template for a query image, clothes regions cropped by a detector from themselves are treated as instruction for gallery images. During the training process, the biometric feature and clothes feature are extracted from person images and instructions. In the inference stage, the model should retrieve images of the same person wearing the provided clothes.
Figure 8: Examples of CTCC-ReID task for training and test. Each identity is further split into query and gallery, where query set consists of query person images and clothes templates as instructions, and gallery set consists of target person images and cropped clothes images as instructions.
C.2 Language-instructed ReID
We provide more examples for LI-ReID as shown in Fig. 8. Similar to CTCC-ReID, the instruction of gallery image is several sentences describing pedestrian attributes. We randomly select the description languages from the corresponding person images in gallery and provide to query images as instruction. The model is required to retrieve images of the same person following the provided sentences.
Figure 9: Examples of LI-ReID task for training and test. Each identity is further split into query and gallery, where query set consists of query person images and language instructions (randomly selected from gallery corresponding to the person), and gallery set consists of target person images and description language for themselves as instructions.
Appendix D Details of OmniReID
In the main text, we briefly introduce the number of images and number of tasks of OmniReID. For the evaluation of OmniReID, we introduce the evaluation scenario and evaluation protocols. In this section, we present detailed information on the training dataset and evaluation dataset and discuss the ethical issues of these datasets.
D.1 Dataset Statistics of OmniReID
OmniReID collects 8 publicly available datasets of 4 existing ReID tasks, including Traditional ReID (Trad-ReID), Clothes-Changing ReID (CC-ReID), Clothes Template Based Clothes-Changing ReID (CTCC-ReID) and Language-Instructed ReID (LI-ReID). As shown in Table 6, we perform two training scenarios based on the built benchmark OmniReID: 1) Single-task Learning (STL): Every dataset is treated as a single task, which is trained and tested individually. 2) Multi-task Learning (MTL): The model is optimized by joint training of all the ReID tasks with all the training datasets. The trained model is then evaluated on different tasks with various datasets. For the Trad-ReID, we select widely-used MSMT17, CUHK03 and Market1501 datasets. For CC-ReID Task, we select widely-used PRCC, LTCC and VC-Clothes, considering the variety of domains. For CTCC-ReID and LI-ReID tasks, we label the COCAS+ Real1, COCAS+ Real2 datasets with language instructions and clothes image instructions for training and test. OmniReID forms a total of 121786 images for training and 166055 images for test, which unites all-purpose person retrieval into one instruct-ReID task.
Table 6: Dataset statistics of OmniReID. We summarize both the training and test set images based on the built benchmark. $\star$ means the images is the same as the LI-ReID task.
Multi-task Single-task Training Test
Dataset Number of images Dataset Number of images
Trad-ReID(MTL) Trad-ReID(STL) MSMT17 [40] 30248 MSMT17 [40] 93820
Trad-ReID(STL) CUHK03 [25] 7365 CUHK03 [25] 6732
Trad-ReID(STL) Market1501 [48] 12936 Market1501 [48] 23100
LI-ReID(STL) LI-ReID(STL) COCAS+ Real1 [23] 34469 COCAS+ Real2 [23] 14449
CC-ReID(MTL) CC-ReID(STL) PRCC [43] 17896 PRCC [43] 10800
CC-ReID(STL) VC-Clothes [39] 9449 VC-Clothes [39] 9611
CC-ReID(STL) LTCC [32] 9423 LTCC [32] 7543
CTCC-ReID(MTL) CTCC-ReID(STL) COCAS+ Real1 [23] $34469^{\star}$ COCAS+ Real2 [23] ${}^{\star}$
Total - - 121786 - 166055
D.2 Discussion of Ethical Issues
The usage of OmniReID might bring several risks, such as privacy, and problematic content. We discuss these risks and their mitigation strategies as follows.
Copyright. All images in this paper and dataset are collected by publicly available. We claim the dataset: (1) Copy and redistribute the material in any medium or format. (2) Remix, transform and build on the material for any purpose, even commercially.